OpenStudy (anonymous):

Find the slope of the line passing through the points (-3,-9) and (7,-9)

OpenStudy (anonymous):

Find the slope of the line passing through the points (-3,-9) and (7,-9)

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

Sameer bhai first find the slope by m = y2 - y1/x2 -x1 then use slope point form with using any point y - y1 = m(x - x1)

OpenStudy (anonymous):

Do you get it ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-9-(-9)=0 7-(-3)=10 0/10=0, so the slope of the line is 0 You could have also seen this right away because the "y" values are the same.

OpenStudy (anonymous):

More proof.